OpenStudy (anonymous):

A family of 3 children and 2 adults visited a health club. The club charges $y an hour for each child and $x an hour for each adult. If the family does not want to spend more than $30 an hour at the health club, graph the best model for this situation. I need to draw a graph for this but im not good at it. can anyone show me how the graph is suppose to look like please?

OpenStudy (anonymous):

@magbak

OpenStudy (anonymous):

Well, they have 3 children at x$ an hour which is the same as 3x, and 2 adults at y$ an hour which is the same as 2y and these added together have to be less than or equal to 30 \(\Large\ 3x+2y\leq 30\) So we can convert this to slope intercept form to easily graph it, subtract 3x \(\Large\ 2y\leq -3x+30\) now divide by 2 \(\Large y\leq \frac{-3}{2} x+30\) To graph this you would put a point at (0,30) then go down 3 and to the right 2 and put a point at (2,27) Then connect the line and shade to the left of the line like that